# Encoding Tactile Stimuli for Organoid Intelligence in Braille Recognition

### v1

@@ Abstract
This study proposes a generalizable encoding strategy that maps tactile sensor data to electrical stimulation patterns, enabling neural organoids to perform an open-loop artificial tactile Braille classification task. Human forebrain organoids cultured on a low-density microelectrode array (MEA) are systematically stimulated to characterize the relationship between electrical stimulation parameters (number of pulse, phase amplitude, phase duration, and trigger delay) and organoid responses, measured as spike activity and spatial displacement of the center of activity. Implemented on event-based tactile inputs recorded from the Evetac sensor, our system achieved an average Braille letter classification accuracy of 61% with a single organoid, which increased significantly to 83% when responses from a three-organoid ensemble were combined. Additionally, the multi-organoid configuration demonstrated enhanced robustness against various types of artificially introduced noise. This research demonstrates the potential of organoids as low-power, adaptive bio-hybrid computational elements and provides a foundational encoding framework for future scalable bio-hybrid computing architectures.
@@ 1 Introduction
We present a universal spatio-temporal-intensity encoding that converts event-based tactile data into current-controlled electrical stimulation for organoids.
@@ 2.1 Organoid and Recording Platform
The platform consists of a remote accessible electrophysiology system integrating four custom air-liquid interface MEAs, each accommodating four organoids with eight electrodes per spheroid for a total of 32 channels per MEA. Voltage signals are acquired via Intan RHS headstages sampling at 30 kHz with 16-bit resolution. All hardware functions are accessible through a Python API and Jupyter notebook interface, with data continuously archived in an InfluxDB time-series database which can be used for downstream analysis.
@@ 2.2 The Impact of Electrode Stimulation Parameters on Organoid Responses
As shown in Figure 2, electrical stimulation on the Neuroplatform is parameterized through several key variables [16]. The “shape” of the current waveform may be biphasic, biphasic with an interphase delay, or triphasic. Each phase is specified by an amplitude and a duration: Amplitude 1 and Duration 1 for the positive phase, Amplitude 2 and Duration 2 for the negative phase. The polarity of the waveform can be inverted to begin with either a negative or a positive phase. Stimulation can be triggered as a single pulse, as a sequence defined in a trigger table, and can be directed to any subset of the 32 electrodes. The trigger delay parameter controls the onset timing of the first pulse, while the number of pulses parameter specifies how many pulses are delivered in each sequence. In all experiments, we chose a positive first and negative second biphasic voltage waveform because it statistically produces the lowest neural activation thresholds while safeguarding both the electrode and surrounding cells over the long term [20]. To improve the lifetime of both the organoid and the electrodes, we set Phase Amplitude 1 = Phase Amplitude 2 and Phase Duration 1 = Phase Duration 2.
@@ 2.2.1 Organoid Responses Under Different Stimulation Parameters
The experiment involves stimulating each electrode within the same organoid respectively, recording data from all electrodes and averaging across them to observe the overall effect of stimulation parameters on the organoid output. Stimuli are delivered with a 2 s inter-stimulus interval to allow sufficient recovery and avoid temporal summation. We record the number of spikes generated across all channels within a 200 ms window following stimulation and compute the mean over 10 trials. We test various pulse counts (0 to 10), phase amplitudes (0 to 20 A), phase durations (0 to 200 s), and trigger delays (0 to 4000 s). A parameter value of 0 indicates spontaneous activity recorded under the unstimulated (spontaneous) condition. During each individual sweep, the remaining parameters are kept at their default values (a single pulse, 4A phase amplitude, and a 100s phase duration).
@@ 2.2.2 Temporal Dynamics of Organoid Responses
Stimulation parameters explored in this experiment are as follows: phase pulse count (1 - 10 pulses), amplitude (2 - 20 A), and phase duration (25 - 200 s). We sweep one parameter at a time through its range while keeping the other two fixed at their default values of a single pulse, 4 A, and 100 s. Each parameter condition is repeated 10 times with a 2 s inter-stimulus interval to allow recovery. We record spike counts from 100 ms before stimulation to 500 ms after stimulation across all channels and average over 10 trials.
@@ 2.2.3 Spatial Dynamics of Organoid Responses
We first recorded the CA during spontaneous activity to establish a baseline. For each stimulation condition, we recorded spike data from all electrodes within a 500 ms post-stimulation window. We systematically tested and recorded the organoid responses by stimulating each electrode with varying numbers of pulses (2, 4, 6, 8, and 10), phase amplitudes (4, 8, 12, 16, and 20 A), and phase durations (50, 100, 150, 200, and 250 s). The default stimulation parameters are set to a pulse count of 5, a phase amplitude of 10 A, and a phase duration of 150 s, corresponding to the midpoints of their respective parameter ranges. Each condition is repeated 100 times with a 500 ms inter-stimulus interval.
@@ 2.3 Applying Organoid Intelligence to Braille Classification
Earlier stimulation tests showed that variations in pulse amplitude, width and count elicit distinct firing patterns in the organoid. Building on these results, we design an encoding pipeline that converts event based tactile signals into electrical stimulation patterns and applies straightforward machine learning classifiers to the resulting organoid activity.
@@ 2.3.1 Experimental Setup and Data Collection
The overall pipeline is as displayed in Figure 3. First, Braille characters are sampled using the Evetac sensor [21]: for each of the 26 Braille letters, the Evetac is lowered to a given depth (0, 0.1, 0.2, 0.3, 0.4 mm) and slid horizontally across the Braille character for 10 trials, yielding 1300 recordings in total. Depth 0 denotes the initial shallow indentation upon contact. Next, we encode this AER data into stimulation patterns for the organoid’s 8 electrodes (as detailed in section 2.3.2) Then stimulation patterns for the different Braille characters are delivered to the organoid and the organoid’s response is read out as an 8-dimensional response vector of spike counts across the organoid’s eight channels. Finally, appropriate classifiers are applied to distinguish among the 26 Braille letters based on organoid responses. We test KNN [24], SVM [18], and random forest [25] classifiers, ultimately selecting SVM as the best-performing option.
@@ 2.3.2 Electrical Encoding of Tactile Event Features
Event count: the total number of events in the region, representing overall tactile intensity.
@@
Time of peak activity: the midpoint time of the 100 ms window with the highest event count, indicating when the Braille character reach that region (original data length was 2000 ms).
@@
We select four key stimulation parameters for the organoids: number of pulses, phase duration, trigger delay, and phase amplitude, and all electrodes use the same encoding range. These parameters control the timing and strength of the stimulation delivered to the organoid. For each AER recording, the features are mapped to stimulation parameters:
@@
We set a 1 s inter-sample interval to allow recovery, and use spikes evoked within 500 ms post-stimulation as organoid output, as shown in Figure 4 for a sample encoding of the letter O.
@@ 2.3.3 Organoid Decoding and Classification Under Noisy Conditions
We decode the spike patterns that emerge from the organoid’s own network dynamics after each stimulus. For every stimulus we count the spikes on each electrode during the 500 ms interval that follows stimulation. Since each organoid is connected to eight electrodes, the resulting spike activity forms an eight-dimensional feature vector, which is used by a support vector machine classifier to identify the Braille character. To study the effect of scale we repeat the experiment with three organoids in parallel. The same eight-region encoding drives each organoid, we record all eight channels from every organoid and concatenate them into a twenty-four-dimensional vector before classification.
Robustness is tested by adding noise to four randomly selected channels, using input data with inherent noise (collected at five different depths). Four noise models are applied: Gaussian noise whose standard deviation equals forty per cent of the channel standard deviation, uniform noise within the same range, data loss that sets forty per cent of the samples to zero, and outliers that multiply forty per cent of the samples by three. Support-vector machine accuracy under each model quantifies how single and triple organoid decoders tolerate interference.
@@ 3.1.1 Organoid Responses Under Different Stimulation Parameters
This experiment tests how spike output varies when one stimulation parameter changes at a time, guiding the choice of safe and informative encoding ranges. Stimulating each electrode in turn and recording from all eight channels reveals clear trends (Figure 5). Spike count increases almost linearly as the number of pulses rises and shows a threshold for phase amplitude: values below 4 A rarely evoke spikes, whereas higher amplitudes produce steadily larger responses. Phase duration produces variable effects and trigger delay has no observable effect on spike count. Baseline recordings without stimulation contain few spikes because spontaneous firing is sparse and the 200 ms observation window is brief.
@@ 3.1.2 Temporal Dynamics of Organoid Responses
This experiment measures how long each stimulus influences network firing. In this experiment, we apply simultaneous stimulation across all electrodes with varying parameters and partition the recording into 50 ms windows, as shown in Figure 6. The red dashed line indicates the 50 ms post-stimulation window midpoint, approximating stimulus time.
When a single pulse is delivered, most responses return to the pre-stimulation activity level within 50 ms. Figure 6(a) shows that increasing pulse count introduces a second peak in spike count, which for 10 pulses can exceed the first peak and extend the response to approximately 100 ms. Figure 6(b) again shows that phase amplitudes below 4 A fail to evoke responses, whereas larger amplitudes do not uniformly increase spike count, different amplitudes produce separable outputs with minimal overlap. Even a 50 µs pulse triggers a clear spike burst (Figure 6(c)), in line with earlier reports [20]. These findings motivate a 500 ms analysis window for each trial in the Braille experiment. The window captures all evoked spikes yet remains short enough that the sparse spontaneous activity in our organoids rarely appears. The Neuroplatform’s adaptive threshold references each electrode’s baseline level, making sure a majority of spikes detected in this window arise from the stimulus.
@@ 3.2 Performance of Organoid in Braille Classification
We evaluate KNN, SVM, and random forest classifiers via cross-validation on 1300 response vectors. SVM achieves the highest accuracy and is used for subsequent analyses. When using only the 8-dimensional output from a single organoid for classification, the accuracy remains stable at approximately 57% (Figure 8(a)). When conducting experiments with three organoids on the same MEA and using their individual 8-dimensional outputs separately for classification, the accuracies are 59%, 62%, and 67%, respectively. However, when combining the outputs of the three organoids into a 24-dimensional vector, the classification accuracy using SVM reaches 83%(Figure 8(b)).
As shown in Figure 8, the heatmaps of organoid responses to each sample show that even though different pressing depths are used during data collection, the response patterns of the organoid remain relatively consistent. which indicates that the organoid network exhibits intrinsic robustness to variations in stimulus strength. Comparing the input and output heatmaps further confirms that a stronger indentation in a given sensor region does not always elicit a stronger response from the electrode assigned to that region. The transformation from tactile input to spike output is therefore shaped by local network dynamics rather than by a direct intensity mapping. When noise is introduced again, the single-organoid system’s accuracy decreases by 2.1–26.5%, with ’missing’ and ’outliers’ noise causing the largest drops; the three-organoid system’s accuracy decreases by 0.8–18.1%. These results demonstrate that organoid redundancy enhances resistance to data disturbances.
@@ 4 Discussion
By systematically characterizing the stimulus-response profile of human forebrain organoids cultured on FinalSpark’s low-density eight-electrode MEA, we developed a unified spatio-temporal-amplitude encoding strategy that converts spike-based outputs from a neuromorphic sensor into electrical-stimulation patterns that can be processed by the organoid network. This transferable approach marks a pivotal step toward leveraging organoid intelligence in open-loop computational tasks. Applied to a blind Braille-letter classification benchmark, the encoder demonstrated competitive performance, and multiplexing three organoids in parallel raised accuracy to 83% while markedly improving robustness against Gaussian, uniform, data loss and outlier noise. Collectively, these results show that even sparsely instrumented organoids can serve as low-power, adaptive classifiers when supplied with appropriately structured input, laying essential groundwork for scalable bio-hybrid computing architectures.
4.1 Toward a Universal Encoding Strategy via Stimulus Parameter Optimization

### v2

# Encoding Tactile Stimuli for Braille Recognition with Organoids

@@ Abstract
This study proposes a transferable encoding strategy that maps tactile sensor data to electrical stimulation patterns, enabling neural organoids to perform an open-loop artificial tactile Braille classification task. Human forebrain organoids cultured on a low-density microelectrode array (MEA) are systematically stimulated to characterize the relationship between electrical stimulation parameters (number of pulse, phase amplitude, phase duration, and trigger delay) and organoid responses, measured as spike activity and spatial displacement of the center of activity. Implemented on event-based tactile inputs recorded from the Evetac sensor, our system achieved an average Braille letter classification accuracy of 61% with a single organoid, which increased significantly to 83% when responses from a three-organoid ensemble were combined. Additionally, the multi-organoid configuration demonstrated enhanced robustness against various types of artificially introduced noise. This research demonstrates the potential of organoids as low-power, adaptive bio-hybrid computational elements and provides a foundational encoding framework for future scalable bio-hybrid computing architectures.
@@ 1 Introduction
We present a spatio-temporal-intensity encoding that converts event-based tactile data into current-controlled electrical stimulation for organoids on low-density MEAs.
@@ 2.1 Organoid and Recording Platform
The platform consists of a remotely accessible electrophysiology system integrating four custom air-liquid interface MEAs, each accommodating four organoids with eight electrodes per organoid for a total of 32 channels per MEA. Voltage signals are acquired via Intan RHS headstages sampling at 30 kHz with 16-bit resolution. All hardware functions are accessible through a Python API and Jupyter notebook interface, with data continuously archived in an InfluxDB time-series database which can be used for downstream analysis. In this study, the Neuroplatform was accessed remotely for stimulation and recording, whereas tactile sensing and data preprocessing were performed locally before stimulation was executed.
@@ 2.2 The Impact of Electrode Stimulation Parameters on Organoid Responses
As shown in Figure 2, electrical stimulation on the Neuroplatform is parameterized through several key variables [21]. The “shape” of the current waveform may be biphasic, biphasic with an interphase delay, or triphasic. A biphasic waveform consists of a positive phase followed by a negative phase, each specified by an amplitude and a duration. A triphasic waveform consists of three phases with alternating polarity. Each phase is specified by an amplitude and a duration: Amplitude 1 and Duration 1 for the positive phase, Amplitude 2 and Duration 2 for the negative phase. The polarity of the waveform can be inverted to begin with either a negative or a positive phase. Stimulation can be triggered as a single pulse, as a sequence defined in a trigger table, and can be directed to any subset of the 32 electrodes. The trigger delay parameter controls the onset timing of the first pulse, while the number of pulses parameter specifies how many pulses are delivered in each sequence. Charge balanced biphasic stimulation is widely used because its primary objective is to keep the electrode potential within a safe range, thereby reducing irreversible electrochemical reactions that can degrade the electrode or damage tissue [10]. In all experiments, we chose a positive first and negative second biphasic voltage waveform because it statistically produces the lowest neural activation thresholds while safeguarding both the electrode and surrounding cells over the long term [38]. To improve the lifetime of both the organoid and the electrodes, we set Phase Amplitude 1 = Phase Amplitude 2 and Phase Duration 1 = Phase Duration 2. We used the Neuroplatform default interphase delay of 0 s between the two phases, and default pulse train period of 10000 s for pulse trains.
@@ 2.2.1 Organoid Responses Under Different Stimulation Parameters
The experiment involves stimulating each electrode within the same organoid respectively, recording data from all electrodes and averaging across them to observe the overall effect of stimulation parameters on the organoid output. Stimuli are delivered with a 2 s inter‐stimulus interval to allow sufficient recovery and avoid temporal summation. Here, one stimulus refers to a single trial consisting of a pulse train with 1 to 10 pulses, the interpulse timing within the train is set by the platform default 10000 s pulse train period. We record the number of spikes generated across all channels within a 200 ms window following stimulation and compute the mean over 10 trials. We test various numbers of pulses (0 to 10), phase amplitudes (0 to 20 A), phase durations (0 to 200 s), and trigger delays (0 to 4000 s). A parameter value of 0 indicates spontaneous activity recorded under the unstimulated (spontaneous) condition. During each individual sweep, the remaining parameters are kept at their default values: a single pulse, 4 A phase amplitude, and a 100 s phase duration. These base parameters were selected to provide a simple and interpretable reference condition for one factor sweeps. We used a single pulse as the minimal stimulation form, minimizing temporal summation effects and prolonged network activation. We set the phase amplitude to 4 A because lower amplitudes were typically subthreshold in our recordings, while higher values risked driving near-saturated responses, and chose a 100 s phase duration as a mid-range value within the sweep to avoid boundary effects.
@@ 2.2.2 Temporal Dynamics of Organoid Responses
Stimulation parameters explored in this experiment are as follows: numbers of pulses (1 - 10 pulses), amplitude (2 - 20 A), and phase duration (25 - 200 s). We sweep one parameter at a time through its range while keeping the other two fixed at their default values of a single pulse, 4 A, and 100 s. Each parameter condition is repeated 10 times with a 2 s inter-stimulus interval to allow recovery. We record spike counts from 100 ms before stimulation to 500 ms after stimulation across all channels and average over 10 trials.
@@ 2.2.3 Spatial Dynamics of Organoid Responses
We first recorded the CA during spontaneous activity to establish a baseline. For each stimulation condition, we recorded spike data from all electrodes within a 500 ms post-stimulation window. We systematically tested and recorded the organoid responses by stimulating each electrode with varying numbers of pulses (2, 4, 6, 8, and 10), phase amplitudes (4, 8, 12, 16, and 20 A), and phase durations (50, 100, 150, 200, and 250 s). The default stimulation parameters are set to a pulse count of 5, a phase amplitude of 10 A, and a phase duration of 150 s, corresponding to the midpoints of their respective parameter ranges. This keeps the reference condition away from both near-threshold and near-saturation regimes, so that changes induced by the swept parameter are more readily observable and less affected by boundary effects. Each condition is repeated 100 times with a 500 ms inter-stimulus interval.
@@ 2.3 Applying Organoid Intelligence to Braille Classification
Earlier stimulation tests showed that variations in pulse amplitude, width and count elicit distinct firing patterns in the organoid (see Section 3.1). Building on these results, we design an encoding pipeline that converts event based tactile signals into electrical stimulation patterns and applies straightforward machine learning classifiers to the resulting organoid activity.
@@ 2.3.1 Experimental Setup and Data Collection
The overall pipeline is as displayed in Figure 3. First, Braille characters are sampled using the Evetac sensor [17]: for each of the 26 Braille letters, the Evetac is lowered to a given depth (0, 0.1, 0.2, 0.3, 0.4 mm) and slid horizontally across the Braille character for 10 trials, yielding 1300 recordings in total. Depth 0 denotes the initial shallow indentation upon contact. Next, we encode this AER data into stimulation patterns for the organoid’s 8 electrodes (as detailed in section 2.3.2) Then stimulation patterns for the different Braille characters are delivered to the organoid and the organoid’s response is read out as an 8‐dimensional response vector of spike counts across the organoid’s eight channels. Finally, appropriate classifiers are applied to distinguish among the 26 Braille letters based on organoid responses. We test KNN [12], SVM [11], and random forest [6] classifiers, ultimately selecting SVM as the best-performing option.
@@ 2.3.2 Electrical Encoding of Tactile Event Features
Event count: the total number of events in the region, representing overall tactile intensity. Here, an “event” refers to a single AER pixel event with an address, polarity, and timestamp. The event count for a region is computed as the total number of raw AER events whose pixel addresses fall within that region, summed over the full 2000 ms tactile recording.
@@
Time of peak activity: the midpoint time of the 100 ms window with the highest event count, indicating when the Braille character reaches that region (original data length was 2000 ms).
@@
We select four key stimulation parameters for the organoids: number of pulses, phase duration, trigger delay, and phase amplitude, and all electrodes use the same encoding range. These parameters control the timing and strength of the stimulation delivered to the organoid. For each AER recording, each extracted feature is linearly mapped to its corresponding stimulation parameter:
@@
We set a 1 s inter‐sample interval to allow recovery, and use spikes evoked within 500 ms post‐stimulation as organoid output, as shown in Figure 4 for a sample encoding of the letter O.
@@ 2.3.3 Organoid Decoding and Classification Under Noisy Conditions
We decode the spike patterns that emerge from the organoid’s own network dynamics after each stimulus. For each stimulus, we count the spikes detected on each electrode during the 500 ms interval following stimulation. Since each organoid is connected to 8 electrodes, the resulting spike activity forms an 8-dimensional feature vector, which is then used by a support vector machine classifier to identify the Braille character. To study the effect of scale, we repeat the experiment with 3 organoids in parallel. The same 8-region encoding drives each organoid, we record all 8 channels from every organoid, and we concatenate them into a 24-dimensional feature vector before classification.
To assess robustness, we introduce artificial disturbances to the spike-count features while keeping the original tactile dataset unchanged. The tactile data are collected at 5 different depths, which already introduces inherent variability, the additional disturbances are applied on top of this. Specifically, we randomly select 4 channels in the feature vector and modify only those values. We apply 4 disturbance models: additive Gaussian noise with a standard deviation equal to 40% of the channel standard deviation, additive uniform noise within the same range, data loss where 40% of the selected entries are set to zero, and outliers where 40% of the selected entries are multiplied by 3. We then report support-vector machine accuracy under each model to quantify how single and triple organoid decoders tolerate channel interference. These models are not intended to cover every realistic sensor artifact, but to provide a basic stress test of whether the performance benefit of using multiple organoids remains under several common types of channel corruption.
@@ 3.1.1 Organoid Responses Under Different Stimulation Parameters
This experiment tests how spike output varies when one stimulation parameter changes at a time, guiding the choice of safe and informative encoding ranges. Stimulating each electrode in turn and recording from all eight channels reveals clear trends (Figure 5). Spike count increases almost linearly as the number of pulses rises and shows a threshold for phase amplitude: values below 4 A rarely evoke spikes, whereas higher amplitudes produce steadily larger responses. In this one-factor sweep, phase duration does not produce a consistent monotonic trend in spike count across the tested conditions, and trigger delay has no observable effect on spike count. Under the no stimulation condition, the spike count measured within the same 200 ms post-trigger window is often close to zero. This reflects the combination of sparse spontaneous firing and the short counting window, rather than an absence of spontaneous activity.
@@ 3.1.2 Temporal Dynamics of Organoid Responses
This experiment measures how long each stimulus influences network firing. In this experiment, we apply simultaneous stimulation across all electrodes with varying parameters and partition the recording into 20 ms windows, as shown in Figure 6. The dashed line marks stimulation onset.
When a single pulse is delivered, most responses return to the pre-stimulation activity level within 60 ms. Figure 6(a) shows that increasing the pulse count effectively prolongs the duration of the evoked activity, extending the response to approximately 100 ms at higher pulse counts. Figure 6(b) again shows that phase amplitudes below 4 A fail to evoke responses, whereas larger amplitudes do not uniformly increase spike count, they result in distinguishable intensities of activation within the post-stimulation window. Even a 50 µs pulse triggers a clear spike burst (Figure 6(c)), in line with earlier reports [38]. These findings motivate a 500 ms analysis window for each trial in the Braille experiment. The window captures all evoked spikes yet remains short enough that the sparse spontaneous activity in our organoids rarely appears. The Neuroplatform’s adaptive threshold references each electrode’s baseline level, making sure a majority of spikes detected in this window arise from the stimulus.
@@ 3.2 Performance of Organoid in Braille Classification
We use a support vector machine (SVM) classifier with cross-validation on 1300 response vectors for all classification analyses reported in this study. When using only the 8-dimensional output from a single organoid for classification, the accuracy remains stable at approximately 57% (Figure 8(a)). When conducting experiments with three organoids on the same MEA and using their individual 8-dimensional outputs separately for classification, the accuracies are 59%, 62%, and 67%, respectively. However, when combining the outputs of the three organoids into a 24-dimensional vector, the classification accuracy using SVM reaches 83% (Figure 8(b)).
As shown in Figure 8, the heatmaps of organoid responses to each sample show that even though different pressing depths are used during data collection, the response patterns of the organoid remain relatively consistent, which indicates that the organoid network exhibits intrinsic robustness to variations in stimulus strength. Comparing the input and output heatmaps further confirms that a stronger indentation in a given sensor region does not always elicit a stronger response from the electrode assigned to that region. The transformation from tactile input to spike output is therefore shaped by local network dynamics rather than by a direct intensity mapping. When noise is introduced again, the single-organoid system’s accuracy decreases by 2.1–26.5% relative to its clean (no-noise) baseline, with ’missing’ and ’outliers’ noise causing the largest drops; the three-organoid system’s accuracy decreases by 0.8–18.1% relative to its clean baseline. Here, the reported ranges are taken across the four noise models. These results demonstrate that organoid redundancy enhances resistance to data disturbances.
@@ 4 Discussion
By systematically characterizing the stimulus-response profile of human forebrain organoids cultured on FinalSpark’s low-density eight-electrode MEA, we developed a transferable spatio-temporal-amplitude encoding strategy that converts spike-based outputs from a neuromorphic sensor into electrical-stimulation patterns that can be processed by the organoid network. Applied to a blind Braille-letter classification benchmark, the encoder demonstrated competitive performance, and multiplexing three organoids in parallel raised accuracy to 83% while markedly improving robustness against Gaussian, uniform, data loss and outlier noise. Collectively, these results show that even sparsely instrumented organoids can support reliable open loop classification when supplied with appropriately structured stimulation input, laying essential groundwork for scalable bio-hybrid computing architectures.
4.1 Toward a Transferable Encoding Strategy via Stimulus Parameter Optimization